# Probing light sterile neutrinos in left-right symmetric models with displaced vertices and neutrinoless double beta decay

### v1

@@ 5.1 Left-right symmetric models
In mLRSM scenarios, determining the precision of the multi-meson corrections is more subtle. In a type-II seesaw scenario without $W_{L}$-$W_{R}$ mixing, all mesons produced in sterile-neutrino decays couple to $W_{R}^{\pm}$ gauge bosons, whereas in the minimal scenario, they couple to $W_{L}^{\pm}$ gauge bosons. Since QCD is invariant under parity transformations, we expect the multi-meson corrections to hold similarly in these scenarios. In this case, employing Eq. (53) would be appropriate for sterile-neutrino masses above 1 GeV. In principle, this is no longer true once we consider non-zero $W_{L}$-$W_{R}$ mixing scenarios since the $\xi$ dependence is dictated by the parity of the hadronic final state [in analogy to Eqs. (50) and (51)]. Because $\xi<0.8$ cannot be too large in practice Maiezza:2010ic , we neglect this difference and use Eq. (53) also when including $W_{L}$-$W_{R}$ mixing. We do not expect this approximation to affect our results significantly but stress that accounting for these $W_{L}$-$W_{R}$ mixing effects by explicitly computing the QCD corrections would be interesting.
@@
Figure 4: Ratios of the sterile-neutrino decay widths into individual mesons to the inclusive hadronic decay width calculated with Eq. (53) for $\xi=0.3$ and $M_{6}=100$, as functions of $M_{4}$/$m_{1}$. The solid line labeled with “$N_{4}\to\text{total}$” corresponds to the sum of the dashed curves.
@@
In Fig. 6, the primary decay branching ratios of the lightest sterile neutrino $N_{4}$ into SM leptons and bound state mesons are depicted for $M_{4}<1.55\,$GeV. The hadronic decays dominate the sterile-neutrino decay processes once they become kinematically accessible around the charged pion threshold $M_{4}>M_{\pi^{\pm}}=139.6\,$MeV. The vector meson $N_{4}\to e^{-}/\mu^{-}+\rho^{+}$ decay modes dominate at larger $M_{4}$, which is unsurprising as a result of the constructive interference terms in the decay rates being linear in $\xi$ and the fact that these processes are Cabibbo-allowed. These branching fractions are independent of the charged RH gauge boson mass $M_{W_{R}}$ since all decay rates scale as $M_{W_{R}}^{-4}$.
@@
In Fig. 7, the proper decay lengths for the lightest sterile neutrino $N_{4}$ as functions of $M_{4}$ are shown in type-I seesaw scenarios without $W_{L}$-$W_{R}$ mixing for various values of the lightest active neutrino’s mass $m_{1}$ and the RH gauge boson’s mass $M_{W_{R}}$. The maximum value $m_{1}=3\times 10^{-2}\,$eV is close to the upper limit permissible under the cosmological constraints on the active neutrino masses. The left (right) panel shows the scenario where $M_{W_{R}}=7\,$TeV (505050\,50TeV). For $M_{W_{R}}=7\,$TeV, when the sterile-neutrino mass surpasses the charged-pion threshold, and RH currents start contributing to the sterile-neutrino decays, the dependence of the sterile-neutrino lifetime on the lightest sterile neutrino mass immediately diminishes. This is due to the relatively light mass of the charged RH gauge boson. For $M_{4}<M_{\pi^{\pm}}$, the only kinematically available decay channels are neutral-current decays, which depend linearly on $m_{1}$. In contrast, the sterile-neutrino lifetimes for $M_{W_{R}}=50\,$TeV exhibit a more substantial dependence on $m_{1}$ across the relevant $M_{4}$ mass range. This dependence becomes progressively less significant as $M_{4}$ increases. The proper-lifetime behavior is similar in a scenario where $\xi=0.3$.
@@ 6 Neutrinoless double beta decay
First of all, with just LH weak interactions, additional contributions arise from the exchange of the heavier sterile neutrinos due to the mixing between $\nu_{L}$ and $\nu_{R}$. Then, in the mLRSM, owing to the RH gauge interactions, there are contributions solely from $\nu_{R}$ exchange. The $W_{L}$-$W_{R}$ mixing also leads to additional topologies. Further, there can be diagrams involving minimal mixing and right-handed gauge interactions (sometimes these are called the $\lambda$ and $\eta$ mechanisms Doi:1985dx ). Finally, there are contributions from the exchange of the doubly charged scalars $\delta_{L,R}^{++}$. However, in scenarios with relatively light RH neutrinos, these contributions are sub-leading and can be neglected deVries:2022nyh . Which of these contributions dominates depends on the interplay between the sterile-neutrino masses, the $M_{W_{L}}/M_{W_{R}}$ mass ratio, and the amount of $W_{L}$-$W_{R}$ mixing.
@@
We recall that in the mLRSM$\big{(}C_{\mathrm{VLL}}^{(6)}\big{)}_{udei}=-2V_{ud}(PU)_{ei}$. In the type-II seesaw limit, the $6\times 6$ $U$-matrix is block-diagonal, and the LH amplitude only sums over the active neutrinos. In this case, we can neglect the neutrino mass dependence of the NMEs and obtain
@@
Because $\xi<0.8$ Maiezza:2010ic cannot be too large, the change in $\mathcal{A}^{\xi=0}$ in the first term is not too severe. The second term, however, can dominate $\mathcal{A}^{\xi}_{R}$. In particular, for $M_{i}\geq 1$ GeV, we can integrate out the sterile neutrinos, and the $W_{L}$-$W_{R}$ mixing leads to dimension-nine lepton-number-violating operators of the form $(\bar{u}_{L}\gamma^{\mu}d_{L})\,(\bar{u}_{R}\gamma^{\mu}d_{R})\,\bar{e}_{L}C\bar{e}_{L}^{T}$. This operator has the right chiral properties to induce, at lower energies, processes of the form $\pi^{-}\pi^{-}\rightarrow e^{-}e^{-}$, which lead to enhanced $0\nu\beta\beta$ amplitudes once the pions are exchanged between nucleons in a nucleus Prezeau:2003xn ; Cirigliano:2018yza ; Nicholson:2018mwc . This enhancement disappears for smaller $M_{i}\leq m_{\pi}$ but is still present for the sterile neutrino masses we are interested in.
@@
Figure 8: $0\nu\beta\beta$ lifetime of Xe136 in the type-II mLRSM in the NH as a function of the lightest (sterile) neutrino mass. We have set the heaviest sterile neutrino mass at 100 GeV and $M_{W_{R}}=10$ TeV and $\xi=0.3$ (blue) and $\xi=0$ (red). The gray regions correspond to the standard mechanism associated with the exchange of light active Majorana neutrinos without right-handed interactions. The right panel is the same as the left but zoomed in on the mass range most relevant for this work.
@@
To illustrate the constraints that can be set from $0\nu\beta\beta$ experiments, we show the $0\nu\beta\beta$ lifetime of 136Xe in Fig. 8 in the type-II limit in the NH as a function of the lightest neutrino mass $m_{1}$ (or equivalently, the lightest sterile neutrino mass $M_{4}$). We have set $M_{6}=100$ GeV and $M_{W_{R}}=10$ TeV for concreteness and considered two cases for the $W_{L}$-$W_{R}$ mixing parameters: $\xi=0$ (red) and $\xi=0.3$ (blue). The gray regions in both figures indicate the lifetime in the case of just three light Majorana neutrinos without contributions from sterile neutrinos or RH currents (we set $M_{W_{R}}\rightarrow\infty$). For $m_{1}<1$ meV, this leads to lifetimes of the order of $10^{29}$ y (the band is caused by marginalizing the unknown Majorana phases), roughly three orders of magnitude longer than the current best limit from KamLAND-Zen, $T^{0\nu}_{1/2}>3.8\cdot 10^{26}$ y KamLAND-Zen:2024eml , shown by the dashed horizontal line. In the mLRSM, the presence of relatively light sterile neutrinos that are charged under $SU(2)_{R}$, as indicated by the blue and red bands, can significantly reduce the lifetimes, reaching lifetimes around $10^{25}$ y for HNLs with masses of $\mathcal{O}(10^{2})$ MeV. We obtain the width of the bands by marginalizing over the unknown Majorana phases and the phases $\alpha$ and $\theta_{L}$. The effect of $W_{L}$-$W_{R}$ mixing can be severe in this mass range because of the chiral enhancement discussed above. In the right panel of Fig. 8, we show the same results but now zoom in on the mass region for $M_{4}$ most relevant for displaced vertex searches. We see that $0\nu\beta\beta$ limits rule out a large part of this parameter space even for a $M_{W_{R}}=10$ TeV mass.
@@ 7 Collider and fixed-target analysis
To evaluate the fraction of $N_{\mathrm{prod}}$ observed through sterile-neutrino decays in the detectors, we apply Monte-Carlo techniques to determine the geometrical acceptance for each detector setup. For a single $N$ with speed $\beta_{i}$, boost factor $\gamma_{i}$ and proper lifetime $\tau_{N}$ (see Sec. 5), the probability to decay inside the fiducial volume of a detector is given by
@@
Here, $L_{T,i}$ is the length $N_{i}$ travels to reach the detector, $L_{I,i}$ is the length $N_{i}$ would travel through the detector if it would not decay inside, and $c$ is the speed of light. In the case of $N_{i}$ traveling not towards the detector, the decay probability inside the detector is zero. The charm and bottom mesons can be legitimately assumed to be decaying at the IP. We cannot do so for the kaons with their longer lifetimes. We also consider their displaced decay positions within the simulation. Only the kaons decaying before reaching areas of rock, shielding infrastructure, hadron absorber, calorimeters, etc., should be included in the computation. For DUNE, there is no such problem as a specific decay pipe is instrumented at the near-detector experiment, while for the far detectors at the LHC, care should be taken and we refer to Ref. Gunther:2023vmz for the detail. For the SHiP experiment, the target is 1.4 m long, followed immediately by a hadron absorber and then muon shielding; thus, we require that the kaons decay before leaving the target and reaching the absorber.
Using the event generator PYTHIA8 Bierlich:2022pfr ; Sjostrand:2014zea in combination with a three-dimensional approximation of the fiducial volume Gunther:2023vmz  [^4] In the three-dimensional approximation we can include a meson traveling a macroscopic distance before decaying into the sterile neutrino. This allows us to include decays of the long-lived kaons in the analysis., we simulate $N_{\mathrm{MC}}=10^{6}$ proton-on-proton events with the corresponding beam-energy setups, and calculate an average decay probability with
$\big{<}P_{M}\big{[}N_{i}\text{ }\mathrm{in}\text{ }\mathrm{f.v.}\big{]}\big{>}=\frac{1}{N_{\mathrm{MC}}}\sum_{i}^{N_{\mathrm{MC}}}P_{M,i}\big{[}N\text{ }\mathrm{in}\text{ }\mathrm{f.v.}\big{]}$ (66)
@@
$N_{N}^{\mathrm{obs}}=\mathrm{Br}\big{(}N\rightarrow\mathrm{visible}\big{)}\sum_{M}N_{M,N}^{\mathrm{prod}}\cdot\big{<}P_{M}\big{[}N\text{ }\mathrm{in}\text{ }\mathrm{f.v.}\big{]}\big{>}\,,$ (67)
where $\mathrm{Br}\big{(}N\rightarrow\mathrm{visible}\big{)}$ is the branching ratio of the sterile neutrino into visible final states, and we assume 100% detection efficiency.
@@ 8.1 Type-II seesaw benchmark scenarios
We first analyze type-II seesaw scenarios with ($\xi=0.3$) and without ($\xi=0.0$) $W_{L}$-$W_{R}$ mixing. In Fig. 10, we have drawn isocurves of $N_{N}^{\text{obs}}=3$ in the $M_{4}\!-\!M_{W_{R}}$ plane for the LHC far detectors, DUNE near detector, and SHiP, as introduced in Sec. 7. To simplify the presentation while obtaining the sensitivity reach to the RH gauge-boson mass, we assume a scenario where $M_{5,6}>m_{B_{s}}$, ensuring that only the lightest sterile neutrino $N_{4}$ is kinematically relevant. We achieve this by selecting a sizeable heaviest sterile-neutrino mass $M_{6}=100\,$GeV, cf. Eq. (32). Note that choosing a substantially larger $M_{6}=1$ TeV would not significantly impact the results. Below the kaon threshold $M_{4}<M_{K^{\pm}}$, DUNE has the highest sensitivity, whereas for $M_{K}^{\pm}<M_{4}<M_{D}^{\pm}$ and $M_{D}^{\pm}<M_{4}<M_{B_{s}}$, SHiP and MATHUSLA are predicted to have the strongest constraining power among the considered experiments, respectively. Furthermore, we display together the latest exclusion bounds of the $0\nu\beta\beta$ experiment KamLAND-Zen KamLAND-Zen:2024eml on the RH gauge-boson mass. These limits were obtained by calculating the ${}^{136}\text{Xe}$ lifetime, as discussed in Sec. 6. Here, to set the most conservative limits on $M_{W_{R}}$, we have maximized the ${}^{136}\text{Xe}$ lifetime over the unknown phases $\lambda_{1,2}$ and $\theta_{L}$. Additionally, we depict the sensitivity reach of next-gen ton-scale experiments that are projected to probe the ${}^{136}\text{Xe}$ lifetime up to $\mathcal{O}(10^{28})$ years.
@@
Figure 10: Predicted sensitivity reach of the next-gen $0\nu\beta\beta$ experiments and the DV searches (at the LHC far detectors, DUNE near detector, and beam-dump experiment SHiP) in the $M_{4}-M_{W_{R}}$ plane in type-II seesaw scenarios. For the $W_{L}$-$W_{R}$ mixing, we have set $\xi=0.3$ ($\xi=0$) in the left (right) panel. Also displayed are existing bounds obtained by recasting the results from the $0\nu\beta\beta$ experiment KamLAND-Zen KamLAND-Zen:2024eml and the search for long-lived sterile neutrinos at the NA62 NA62:2020mcv , T2K T2K:2019jwa , and BEBC WA66:1985mfx ; Barouki:2022bkt experiments. We have set the heaviest sterile neutrino mass $M_{6}=100$ GeV.
Below the kaon threshold, the most stringent upper limits on the minimal mixing angle $|U_{e4}|^{2}$ between the active and sterile neutrinos have been determined by the experiments NA62 NA62:2020mcv and T2K T2K:2019jwa . We translate these upper bounds into lower bounds on the RH gauge-boson mass, which agree with previous assessments Alves:2023znq . We can compare these limits to the $0\nu\beta\beta$ constraints. For $\xi=0.0$, current NA62 and T2K limits are competitive with current $0\nu\beta\beta$ constraints, whereas, for larger $\xi$, the $0\nu\beta\beta$ limits are more stringent. Upper bounds on the minimal mixing angle have also been determined by experiments such as CHARM CHARM:1985nku ; Orloff:2002de , DELPHI DELPHI:1996qcc , PS191 Bernardi:1987ek ; Ruchayskiy:2011aa , BEBC WA66:1985mfx ; Barouki:2022bkt , and NuTeV NuTeV:1999kej . Below the kaon threshold, these experiments’ constraints are weaker than those of NA62 NA62:2020mcv and T2K T2K:2019jwa , and we therefore do not show them. However, BEBC WA66:1985mfx ; Barouki:2022bkt does extend the limits to larger sterile-neutrino masses leading the bounds on $|U_{e4}|^{2}$ for $M_{4}$ between the kaon and charm-meson thresholds, and its bounds are shown as dashed light-gray curves. They are much less stringent than the $0\nu\beta\beta$ limits.

### v2

@@ 5.1 Left-right symmetric models
In mLRSM scenarios, determining the precision of the multi-meson corrections is more subtle. In a type-II seesaw scenario without $W_{L}$-$W_{R}$ mixing, all mesons produced in sterile-neutrino decays couple to $W_{R}^{\pm}$ gauge bosons, whereas in the minimal scenario, they couple to $W_{L}^{\pm}$ gauge bosons. Since QCD is invariant under parity transformations, we expect the multi-meson corrections to hold similarly in these scenarios. In this case, employing Eq. (53) would be appropriate for sterile-neutrino masses above 1 GeV. In principle, this is no longer true once we consider non-zero $W_{L}$-$W_{R}$ mixing scenarios since the $\xi$ dependence is dictated by the parity of the hadronic final state [in analogy to Eqs. (50) and (51)]. Because $\xi<0.8$ cannot be too large in practice Maiezza:2010ic , we neglect this difference and use Eq. (53) also when including $W_{L}$-$W_{R}$ mixing. We do not expect this approximation to significantly affect our results, but we stress that explicitly computing the QCD corrections would be interesting for accounting for these $W_{L}$-$W_{R}$ mixing effects.
@@
Figure 4: Ratios of the sterile-neutrino decay widths into individual mesons to the inclusive hadronic decay width calculated with Eq. (53) for $\xi=0.3$ and $M_{6}=100$ GeV, as functions of $M_{4}$/$m_{1}$. The solid line labeled with “$N_{4}\to\text{total}$” corresponds to the sum of the dashed curves.
@@
In Fig. 6, the primary decay branching ratios of the lightest sterile neutrino $N_{4}$ into SM leptons and bound-state mesons are depicted for $M_{4}<1.55\,$GeV. The hadronic decays dominate the sterile-neutrino decay processes once they become kinematically accessible around the charged pion threshold $M_{4}>M_{\pi^{\pm}}=139.6\,$MeV. The vector meson $N_{4}\to e^{-}/\mu^{-}+\rho^{+}$ decay modes dominate at larger $M_{4}$, which is unsurprising as a result of the constructive interference terms in the decay rates being linear in $\xi$ and the fact that these processes are Cabibbo-allowed. These branching fractions are independent of the charged RH gauge-boson mass $M_{W_{R}}$ since all decay rates scale as $M_{W_{R}}^{-4}$.
@@
In Fig. 7, the proper decay lengths for the lightest sterile neutrino $N_{4}$ as functions of $M_{4}$ are shown in type-I seesaw scenarios without $W_{L}$-$W_{R}$ mixing for various values of the lightest active neutrino’s mass $m_{1}$ and the RH gauge boson’s mass $M_{W_{R}}$. The maximum value $m_{1}=3\times 10^{-2}\,$eV is close to the upper limit permissible under the cosmological constraints on the active-neutrino masses Giusarma:2016phn ; Vagnozzi:2017ovm ; Vagnozzi:2018jhn ; Tanseri:2022zfe ; Planck:2018vyg . The left (right) panel shows the scenario where $M_{W_{R}}=7\,$TeV (505050\,50TeV). For $M_{W_{R}}=7\,$TeV, when the sterile-neutrino mass surpasses the charged-pion threshold, and RH currents start contributing to the sterile-neutrino decays, the dependence of the sterile-neutrino lifetime on the lightest sterile-neutrino mass immediately diminishes. This is due to the relatively light mass of the charged RH gauge boson. For $M_{4}<M_{\pi^{\pm}}$, the only kinematically available decay channels are neutral-current decays, which depend linearly on $m_{1}$. In contrast, the sterile-neutrino lifetimes for $M_{W_{R}}=50\,$TeV exhibit a more substantial dependence on $m_{1}$ across the relevant $M_{4}$ mass range. This dependence becomes progressively less significant as $M_{4}$ increases. The proper-lifetime behavior is similar in a scenario where $\xi=0.3$.
@@ 6 Neutrinoless double beta decay
First of all, with just LH weak interactions, additional contributions arise from the exchange of the heavier sterile neutrinos as a result of the mixing between $\nu_{L}$ and $\nu_{R}$. Then, in the mLRSM, owing to the RH gauge interactions, there are contributions solely from $\nu_{R}$ exchange. The $W_{L}$-$W_{R}$ mixing also leads to additional topologies. Further, there can be diagrams involving minimal mixing and right-handed gauge interactions (sometimes these are called the $\lambda$ and $\eta$ mechanisms Doi:1985dx ). Finally, there are contributions from the exchange of the doubly charged scalars $\delta_{L,R}^{++}$. However, in scenarios with relatively light RH neutrinos, these contributions are sub-leading and can be neglected deVries:2022nyh . Which of these contributions dominates depends on the interplay between the sterile-neutrino masses, the $M_{W_{L}}/M_{W_{R}}$ mass ratio, and the amount of $W_{L}$-$W_{R}$ mixing.
@@
We recall that in the mLRSM$\big{(}C_{\mathrm{VLL}}^{(6)}\big{)}_{udei}=-2V_{ud}(PU)_{ei}$. In the type-II seesaw limit, the $6\times 6$ $U$-matrix is block-diagonal, and the LH amplitude only sums over the active neutrinos. In this case, we can neglect the neutrino-mass dependence of the NMEs and obtain
@@
Because $\xi<0.8$ Maiezza:2010ic cannot be too large, the change in $\mathcal{A}^{\xi=0}$ in the first term is not too severe. The second term, however, can dominate $\mathcal{A}^{\xi}_{R}$. In particular, for $M_{i}\geq 1$ GeV, we can integrate out the sterile neutrinos, and the $W_{L}$-$W_{R}$ mixing leads to dimension-nine lepton-number-violating operators of the form $(\bar{u}_{L}\gamma^{\mu}d_{L})\,(\bar{u}_{R}\gamma^{\mu}d_{R})\,\bar{e}_{L}C\bar{e}_{L}^{T}$. This operator has the right chiral properties to induce, at lower energies, processes of the form $\pi^{-}\pi^{-}\rightarrow e^{-}e^{-}$, which lead to enhanced $0\nu\beta\beta$ amplitudes once the pions are exchanged between nucleons in a nucleus Prezeau:2003xn ; Cirigliano:2018yza ; Nicholson:2018mwc . This enhancement disappears for smaller $M_{i}\leq m_{\pi}$ but is still present for the sterile-neutrino masses we are interested in.
@@
Figure 8: $0\nu\beta\beta$ lifetime of Xe136 in the type-II mLRSM in the NH as a function of the lightest (sterile) neutrino mass. We have set the heaviest sterile-neutrino mass at 100 GeV and $M_{W_{R}}=10$ TeV and $\xi=0.3$ (blue) and $\xi=0$ (red). The gray regions correspond to the standard mechanism associated with the exchange of light active Majorana neutrinos without right-handed interactions. The right panel is the same as the left but zoomed in on the mass range most relevant for this work.
@@
To illustrate the constraints that can be set from $0\nu\beta\beta$ experiments, we show the $0\nu\beta\beta$ lifetime of 136Xe in Fig. 8 in the type-II limit in the NH as a function of the lightest neutrino mass $m_{1}$ (or equivalently, the lightest sterile-neutrino mass $M_{4}$). We have set $M_{6}=100$ GeV and $M_{W_{R}}=10$ TeV for concreteness and considered two cases for the $W_{L}$-$W_{R}$ mixing parameters: $\xi=0$ (red) and $\xi=0.3$ (blue). The gray regions in both figures indicate the lifetime in the case of just three light Majorana neutrinos without contributions from sterile neutrinos or RH currents (we set $M_{W_{R}}\rightarrow\infty$). For $m_{1}<1$ meV, this leads to lifetimes of the order of $10^{29}$ y (the band is caused by marginalizing the unknown Majorana phases), roughly three orders of magnitude longer than the current best limit from KamLAND-Zen, $T^{0\nu}_{1/2}>3.8\cdot 10^{26}$ y KamLAND-Zen:2024eml , shown by the dashed horizontal line. We note that even the next-generation experiments, nEXO and LEGEND, projected to improve the sensitivities by a factor of $\sim 100$ nEXO:2017nam ; LEGEND:2017cdu , cannot reach that, as shown also with a dashed horizontal line. In the mLRSM, the presence of relatively light sterile neutrinos that are charged under $SU(2)_{R}$, as indicated by the blue and red bands, can significantly reduce the lifetimes, reaching lifetimes around $10^{25}$ y for HNLs with masses of $\mathcal{O}(10^{2})$ MeV. We obtain the width of the bands by marginalizing over the unknown Majorana phases and the phases $\alpha$ and $\theta_{L}$. The effect of $W_{L}$-$W_{R}$ mixing can be severe in this mass range because of the chiral enhancement discussed above. In the right panel of Fig. 8, we show the same results but now zoom in on the mass region for $M_{4}$ most relevant for displaced-vertex searches. We see that $0\nu\beta\beta$ limits rule out a large part of this parameter space even for a $M_{W_{R}}=10$ TeV mass.
@@ 7 Collider and fixed-target analysis
To evaluate the fraction of $N_{\mathrm{prod}}$ observed through sterile-neutrino decays in the detectors, we apply Monte-Carlo techniques to determine the geometrical acceptance for each detector setup. For a single $N$ with speed $\beta_{i}$, boost factor $\gamma_{i}$ and proper lifetime $\tau_{N}$ (see Sec. 5) produced in the decay of a meson $M$, the probability of decaying inside the fiducial volume of a detector is given by
@@
Here, $L_{T,i}$ is the length $N_{i}$ travels to reach the detector, $L_{I,i}$ is the length $N_{i}$ would travel through the detector if it would not decay inside, and $c$ is the speed of light. The values of $L_{T,i}$ and $L_{I,i}$ depend not only on the experiment on hand but also on the traveling direction of $N_{i}$ as well as the vertex at which the meson decays into $N_{i}$. To understand their typical values, we refer to Refs. DeVries:2020jbs ; Gunther:2023vmz for a summary of the geometrical setup of the experiments considered in this work. If $N_{i}$ does not travel towards the detector, the decay probability inside the detector is zero. The charm and bottom mesons can be legitimately assumed to be decaying at the IP. However, we cannot do so for the kaons with their longer lifetimes. We thus also consider their displaced decay positions within the simulation. The computation should include only the kaons decaying before reaching areas of rock, shielding infrastructure, hadron absorbers, calorimeters, etc. For DUNE, there is no such problem as a specific decay pipe is instrumented at the near-detector experiment, while for the far detectors at the LHC, care should be taken and we refer to Ref. Gunther:2023vmz for the detail. For the SHiP experiment, the target is 1.4 m long SHiP:2021nfo , followed immediately by a hadron absorber and then muon shielding; thus, we require that the kaons decay before traveling one nuclear interaction length ($\sim 20$ cm) into the target.
Using the event generator PYTHIA8 Bierlich:2022pfr ; Sjostrand:2014zea in combination with a three-dimensional approximation of the fiducial volume Gunther:2023vmz  [^5] In the three-dimensional approximation we can include a meson traveling a macroscopic distance before decaying into the sterile neutrino. This allows us to include decays of the long-lived kaons in the analysis., we simulate $N_{\mathrm{MC}}=10^{6}$ proton-on-proton events with the corresponding beam-energy setups and calculate an average decay probability with
$\big{<}P_{M}\big{[}N_{i}\text{ }\mathrm{in}\text{ }\mathrm{f.v.}\big{]}\big{>}=\frac{1}{N_{\mathrm{MC}}}\sum_{i}^{N_{\mathrm{MC}}}P_{M,i}\big{[}N\text{ }\mathrm{in}\text{ }\mathrm{f.v.}\big{]}.$ (66)
@@
$N_{N}^{\mathrm{obs}}=\mathrm{Br}\big{(}N\rightarrow\mathrm{visible}\big{)}\sum_{M}N_{M,N}^{\mathrm{prod}}\cdot\big{<}P_{M}\big{[}N\text{ }\mathrm{in}\text{ }\mathrm{f.v.}\big{]}\big{>}\cdot h_{M}^{\text{decay}}\,,$ (67)
where $\mathrm{Br}\big{(}N\rightarrow\mathrm{visible}\big{)}$ is the branching ratio of the sterile neutrino into visible final states, and we assume 100% detection efficiency. $h_{M}^{\text{decay}}$ is the fraction of the mesons $M$ decaying in time, cf. the discussion below Eq. (65). It is, therefore, equal to one for the charm and bottom mesons but is non-zero for kaons at the LHC far detectors and the SHiP experiment.
@@ 8.1 Type-II seesaw benchmark scenarios
We first analyze type-II seesaw scenarios with ($\xi=0.3$) and without ($\xi=0.0$) $W_{L}$-$W_{R}$ mixing. In Fig. 10, we have drawn isocurves of $N_{N}^{\text{obs}}=3$ in the $M_{4}\!-\!M_{W_{R}}$ plane for the LHC far detectors, DUNE near detector, and SHiP, as introduced in Sec. 7. To simplify the presentation while obtaining the sensitivity reach to the RH gauge-boson mass, we assume a scenario where $M_{5,6}>m_{B_{s}}$, ensuring that only the lightest sterile neutrino $N_{4}$ is kinematically relevant. We achieve this by selecting a sizable heaviest sterile-neutrino mass $M_{6}=100\,$GeV, cf. Eq. (32). Choosing a substantially larger $M_{6}=1$ TeV would not significantly impact the results. Below the kaon threshold $M_{4}<M_{K^{\pm}}$, DUNE has the highest sensitivity, whereas for $M_{K}^{\pm}<M_{4}<M_{D}^{\pm}$ and $M_{D}^{\pm}<M_{4}<M_{B_{s}}$, SHiP and MATHUSLA are predicted to have the strongest constraining power among the considered experiments, respectively. Furthermore, we display together the latest exclusion bounds of the $0\nu\beta\beta$ experiment KamLAND-Zen KamLAND-Zen:2024eml on the RH gauge-boson mass. These limits were obtained by calculating the ${}^{136}\text{Xe}$ lifetime, as discussed in Sec. 6. Here, to set the most conservative limits on $M_{W_{R}}$, we have maximized the ${}^{136}\text{Xe}$ lifetime over the unknown phases $\lambda_{1,2}$ and $\theta_{L}$. Additionally, we depict the sensitivity reach of next-gen ton-scale experiments that are projected to probe the ${}^{136}\text{Xe}$ lifetime up to $\mathcal{O}(10^{28})$ years nEXO:2017nam ; LEGEND:2017cdu . We note that for $M_{4}$ below the pion threshold, the strongest sensitivity reach should, in principle, be derived by considering pion decays. We do not consider these decays in our simulation as our tool PYTHIA8 is not well validated for forward pion production. Therefore, we shade the corresponding parameter region with a light red color. Nevertheless, in this mass regime, the leading bounds on $|U_{e4}|^{2}$ were reported by the PIENU collaboration PIENU:2019usb . We have reinterpreted these bounds in terms of our model and will display them in the numerical results.
@@
Figure 10: Predicted sensitivity reach of the next-gen $0\nu\beta\beta$ experiments and the DV searches (at the LHC far detectors, DUNE near detector, and beam-dump experiment SHiP) in the $M_{4}-M_{W_{R}}$ plane in type-II seesaw scenarios. For the $W_{L}$-$W_{R}$ mixing, we have set $\xi=0.3$ ($\xi=0$) in the left (right) panel. Also displayed are existing bounds obtained by recasting the results from the $0\nu\beta\beta$ experiment KamLAND-Zen KamLAND-Zen:2024eml and the search for long-lived sterile neutrinos at the PIENU PIENU:2019usb , NA62 NA62:2020mcv , T2K T2K:2019jwa , and BEBC WA66:1985mfx ; Barouki:2022bkt experiments. The region for $M_{4}<m_{\pi}$ is shaded in light red, as we do not include sterile neutrinos from pion decays in our simulation. We have set the heaviest sterile neutrino mass $M_{6}=100$ GeV.
Below the kaon threshold, the most stringent upper limits on the minimal mixing angle $|U_{e4}|^{2}$ between the active and sterile neutrinos have been determined by the experiments NA62 NA62:2020mcv and T2K T2K:2019jwa . [^6] As mentioned above, in our numerical simulation we do not take into account the sterile neutrinos from pion decays which should dominate the sensitivity reach for $M_{4}<m_{\pi}$, where the strongest present bounds on $|U_{e4}|^{2}$ were obtained at the PIENU experiment PIENU:2019usb . These bounds have been reinterpreted here in terms of our scenario. We translate these upper bounds into lower bounds on the RH gauge-boson mass, which agree with previous assessments Alves:2023znq . We can compare these limits to the $0\nu\beta\beta$ constraints. For $\xi=0.0$, current NA62 and T2K limits are competitive with current $0\nu\beta\beta$ constraints, whereas, for larger $\xi$, the $0\nu\beta\beta$ limits are more stringent. Upper bounds on the minimal mixing angle have also been determined by experiments such as CHARM CHARM:1985nku ; Orloff:2002de , DELPHI DELPHI:1996qcc , PS191 Bernardi:1987ek ; Ruchayskiy:2011aa , BEBC WA66:1985mfx ; Barouki:2022bkt , and NuTeV NuTeV:1999kej . Below the kaon threshold, these experiments’ constraints are weaker than those of NA62 NA62:2020mcv and T2K T2K:2019jwa , and we therefore do not show them. However, BEBC WA66:1985mfx ; Barouki:2022bkt does extend the limits to larger sterile-neutrino masses leading the bounds on $|U_{e4}|^{2}$ for $M_{4}$ between the kaon and charm-meson thresholds, and its bounds are shown as dashed light-gray curves. They are much less stringent than the $0\nu\beta\beta$ limits.